OpenStudy (anonymous):

If the cost of corn varies directly with the number of ears purchased, and 6 ears cost 9 dollars, how many ears would cost 162 dollars. A. 18 B. 243 C. 486 D. 1/3 E. none of these. <--- answer i'm getting is E the correct answer?

OpenStudy (anonymous):

ear dollar 6 9 x 162 cross multiply

OpenStudy (anonymous):

that would be 162 x 6 = 9x

OpenStudy (anonymous):

i mean\[162 \times 6 = 9x\]

OpenStudy (anonymous):

okay and x would be 108?

OpenStudy (anonymous):

that's what i got when i did the math, so E is correct then?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

anymore questions?

OpenStudy (anonymous):

I made a mistake sorry :)

OpenStudy (anonymous):

Yes @ChiefArnav one more! The roses exceeded the violets by 222. if the number of roses was 60 more than twice the number of violets, how many violets were there? A. 81 B. 54 C. 162 D. 222 E. none of these.

OpenStudy (anonymous):

which i also got E for.

OpenStudy (anonymous):

i got 384 more

OpenStudy (anonymous):

1. r = # of roses and v = # of violets r = v + 222 r = 2v + 60 v + 222 =2v + 60 v = 162, r = 384

OpenStudy (anonymous):

how did you get E?

OpenStudy (anonymous):

anymore questions?

OpenStudy (anonymous):

oh i see what i did. i stopped once i got 384.

OpenStudy (anonymous):

i didnt finish solving.

OpenStudy (anonymous):

nope i think thats all. thank you so much :)

OpenStudy (anonymous):

no problem :)